P and Q represent numbers, and P*Q means (P+Q)/2. What is the value of 3*(6*8)?
the problem is i cant tell what * that star looking thing mean
3*(6*8) = P*Q i think so and
3+(6*8)= P+Q i also think so
3*(6*8)= 3+(6*8)
that 3+(6*8) is to all be divide by 2 but i dont know what * means its not multiplying but a number i am pretty positive on it

Answer by stanbon(75887)

P and Q represent numbers, and P*Q means (P+Q)/2.
Comment:: They are defining * to mean "add P and Q; then divide by 2".
--------
What is the value of 3*(6*8)?
----
Evaluate the parentheses to get:
3*((6+8)/2) = 3*(7)
-----
Ans: 3*7 = (3+7)/2 = 5
--------------------
Cheers,
Stan H.
-----------